### v3

# Without Paired Labeled Data: End-to-End Self-Supervised Learning for Drone-view Geo-Localization

Abstract
Drone-view Geo-Localization (DVGL) aims to achieve accurate localization of drones by retrieving the most relevant GPS-tagged satellite images. However, most existing methods heavily rely on strictly pre-paired drone-satellite images for supervised learning. When the target region shifts, new paired samples are typically required to adapt to the distribution changes. The high cost of annotation and the limited transferability of these methods significantly hinder the practical deployment of DVGL in open-world scenarios. To address these limitations, we propose an end-to-end self-supervised learning method with a shallow backbone network. It employs a clustering algorithm to generate pseudo-labels and adopts a dual-path contrastive learning framework to learn discriminative intra-view representations. Furthermore, our method incorporates two core modules, including the dynamic hierarchical memory learning module (DHML) and the information consistency evolution learning module (ICEL). The DHML combines short-term and long-term memory to enhance intra-view feature consistency and discriminability. Meanwhile, the ICEL module utilizes a neighborhood-driven dynamic constraint mechanism to systematically capture implicit cross-view semantic correlations, consequently improving cross-view feature alignment. To further stabilize and strengthen the self-supervised training process, a pseudo-label enhancement strategy is introduced to enhance the quality of pseudo supervision. Extensive experiments on three public benchmark datasets demonstrate that the proposed method consistently outperforms existing self-supervised methods and even surpasses several state-of-the-art supervised methods. Our code is available at https://github.com/ISChenawei/DMNIL.
Index Terms:
Drone-view Geo-Localization, Self-Supervised Learning, Contrastive Learning.
I Introduction
Drone-View Geo-Localization (DVGL) is intended to achieve high-precision localization of drones in GPS-denied scenarios by retrieving GPS-tagged satellite images that correspond to drone aerial images [1]. This capability is critical for applications such as autonomous navigation, disaster rescue, and emergency response. In recent years, existing DVGL methods have mainly relied on supervised learning using strictly paired drone-satellite images [2, 3]. Despite the promising performance demonstrated by these supervised methods, several intrinsic limitations are exposed. Firstly, the acquisition of instance-level paired drone-satellite images incurs additional training expenses [4]. Secondly, large amounts of unpaired cross-view image resources remain under-exploited. Lastly, when these methods are applied to new scenarios, the need arises for re-collection and re-pairing of drone-satellite images for training purposes to accommodate domain shifts. These limitations significantly hinder the practical deployment of DVGL in open-world scenarios. Thereby, it is necessary to explore a self-supervised method that can extract latent cross-view correlations from large-scale unpaired drone-satellite images.
Figure 1: As representative cross-view retrieval tasks, ReID and DVGL share the common objective of establishing correspondences between query and gallery images based on discriminative feature representations. The notable progress achieved by ReID, especially in self-supervised learning scenarios, offers a valuable methodological foundation and insightful paradigms for advancing DVGL research. However, DVGL presents greater challenges for effective feature learning and discrimination compared with ReID. (A) ReID tasks are characterized by relatively small intra-class variations and distinct inter-class differences, which facilitates effective feature discrimination. (B) In contrast, DVGL tasks suffer from large intra-class variations and small inter-class differences, leading to significant cross-view gaps and intra-view ambiguities that hinder feature discrimination.
In recent years, self-supervised person re-identification (ReID) methods [5, 6] have achieved remarkable progress to offer crucial theoretical foundations for DVGL. Typically, these methods employ clustering algorithms [7, 8] on features extracted by a backbone network to generate pseudo-labels, which serve as the initial supervision signals for self-supervised learning. However, as illustrated in Fig. 1, compared to conventional ReID tasks, DVGL presents more severe challenges, such as large intra-view ambiguities and cross-view discrepancies. As shown in Fig. 2, these issues hinder the backbone network from extracting stable and discriminative features at the early stage, which directly undermines the quality of pseudo-labels generated during the initial clustering phase. Without effective strategies to guide the learning of intra-view and cross-view consistency as well as feature discrimination, the model lacks the necessary ability to extract meaningful representations from low-quality pseudo-labels. As a result, the learning process may precipitate a detrimental feedback loop, wherein inaccurate pseudo-labels undermine the quality of feature representations to further compromise pseudo-label quality in subsequent iterations, ultimately leading to the collapse of self-supervised training. Although a current self-supervised method [9] proposed a coarse-to-fine framework for DVGL with a similar dual-path contrastive learning framework, it relies on explicit appearance-level transformations rather than intrinsically modeling the underlying cross-view feature relationships. In the shared space, it still exhibits limited cross-view alignment and susceptibility to feature noise.
Figure 2: Comparison of Recall@k performance on the DVGL and ReID datasets using ConvNeXt-Tiny [10], which is pre-trained on ImageNet-22k [11]. The evaluation is conducted on the DVGL dataset University-1652 [12] and the four person ReID datasets including MSMT17 [13], DukeMTMC-reID [14], CUHK03 [15], and Market1501 [16].
In this paper, we propose a novel end-to-end self-supervised learning method for DVGL. A lightweight and shallow ConvNeXt-Tiny backbone network [10] is employed to extract features, which are subsequently clustered to generate pseudo-labels. Combined with a dual-path contrastive learning strategy [17, 18], the backbone network coupled with a downstream clustering process forms the baseline network of our method for learning intra-view feature representations. Building upon this baseline, we design the dynamic hierarchical memory learning (DHML) module and the information consistency evolution learning (ICEL) module as two key components of our method. The DHML module captures the relationship between short-term and long-term memory to enhance the discriminability of intra-view feature representations. The ICEL module constructs a neighborhood-driven dynamic constraint mechanism by integrating neighborhood consistency constraints with mutual information optimization to facilitate cross-view feature alignment and discriminative representation learning. DHML enhances intra-view discrimination and ICEL improves cross-view consistency. Through joint optimization of intra-view and cross-view features, they collectively strengthen the feature discriminability and consistency. Additionally, a pseudo-label enhancement (PLE) strategy is employed to improve the reliability of pseudo-labels and guide more stable self-supervised feature learning. The main contributions can be summarized as follows.
  • We propose an end-to-end self-supervised learning method for DVGL. Built upon a shallow backbone and a dual-path contrastive learning strategy, our method integrates dynamic hierarchical memory learning and information consistency evolution learning to enhance feature consistency and discriminability for both intra-view and cross-view representation learning. Additionally, a pseudo-label enhancement (PLE) strategy is employed to guide more stable self-supervised learning.
  • To address the intra-view ambiguity caused by spatial continuity across geographic viewpoints, we propose a dynamic memory hierarchical learning (DHML) module that captures the dependencies between different memory levels to adaptively learn the dynamics of intra-class and inter-class features within the intra-view.
  • To address cross-view discrepancies and inconsistencies, we propose a novel information consistency evolution learning (ICEL) module which combines neighborhood consistency and mutual information optimization to learn consistent and discriminative cross-view representations, bridging the discriminability in intra-view learned by the DHML module to form a synergistic learning system.
  • Extensive experiments demonstrate that our method, which employs a lightweight backbone, outperforms existing self-supervised methods on three DVGL benchmarks and even surpasses several state-of-the-art supervised methods.
The remainder of this paper is organized as follows. Section II systematically reviews previous research. In Section III, the proposed self-supervised learning method is presented in detail. The experimental results are reported and analyzed in Section IV. Finally, conclusions are outlined in Section V.
II Related Works
This section establishes a systematic and comprehensive theoretical framework to support the development of the proposed self-supervised learning method for DVGL. DVGL is typically regarded as a sub-task of cross-view geo-localization (CVGL). To provide a more comprehensive understanding of the field, we first review the overall progress in CVGL, including both supervised and self-supervised methods. Then, we summarize the main methods in self-supervised ReID that have provided analogous problem-solving paradigms.
II-A Supervised Cross-View Geo-localization
CVGL aims to achieve high-precision localization of query images by retrieving GPS-tagged satellite images. Early research focused mainly on the CVGL in ground-view [19] with several benchmark datasets, such as CVUSA [20], CVACT [21], and VIGOR [22]. To address the challenge of ground-view CVGL, researchers have proposed various methods. CVM-Net [23] employed NetVLAD [24] to encode images as global descriptors to reduce visual discrepancies caused by viewpoint variations. TransGeo [25] utilized attention-based zooming and sharpness-aware optimization to enhance detail learning and improve model generalization. GeoDTR [26] leveraged a Transformer-based [27] feature extractor to disentangle geometric features, effectively addressing perspective shifts. Sample4Geo [28] introduced a hard negative mining strategy, focusing on challenging negative samples to further improve discriminative capability.
With the continuous development and application of drone technology, DVGL has gradually become an important research direction and has achieved significant progress [29, 30, 31]. Several benchmark datasets for DVGL have been proposed, such as University-1652 [12], SUES-200 [32], and DenseUAV [33], which have driven further research in this field. In this scenario, researchers have proposed a variety of methods. LPN [34] extracted fine-grained features through local pattern partitioning to improve the ability to capture local details. IFS [35] utilized a multi-branch strategy to effectively fuse global and local features. CAMP [29] improved the accuracy of cross-view matching by contrastive attribute mining and position-aware partitioning, while DAC [3] enhanced feature consistency and matching accuracy by introducing domain alignment and scene consistency constraints. MEAN [2] adopted progressive embedding diversification, global-to-local associations, and cross-domain enhanced alignment to further improve the representation and alignment of features.
Nevertheless, the remarkable performances exhibited by these DVGL methods require extensive pre-processed drone-satellite image pairs. In this work, we concentrate on self-supervised learning approaches for DVGL, particularly addressing scenarios where labeled pairings are unavailable, which is crucial for real-world applications.
II-B Self-Supervised Cross-View Geo-localization
Due to the dependence of CVGL on extensive pre-processed image pairs, which significantly increases costs and limits applications for the real world, researchers have begun to explore self-supervised learning methods for CVGL. The work in [36] proposed a self-supervised learning method that used correspondence-free projection to transform ground panorama images into bird’s eye view images. It also applied CycleGAN [37] to generate fake satellite images, coupled with an additional SAFA module [38] to reduce discrepancies. However, this explicit alignment method has inherent limitations in capturing the deep semantic relationships of cross-view features and may introduce noise. Furthermore, the non-end-to-end training method employed significantly increases training time and computational costs with low practical applicability. Additionally, existing work [4] utilized frozen foundation models (FMs) [39] to extract features, and a self-supervised adapter was introduced to mitigate visual discrepancies between different points of view. However, this method heavily relied on the ability of FMs for feature extraction. When FMs fail to effectively handle viewpoint discrepancies, the performance of the adapter is also affected. To reduce reliance on the feature extractor, CDIKTNet [40] trained with a small amount of paired drone-satellite images to optimize the initial feature distribution and then performed clustering to generate high-quality pseudo-labels. However, this method inherently depends on the feature extractor, as it still requires limited paired drone-satellite images for effective learning.
The above methods have made meaningful exploratory attempts. However, these methods still face challenges, such as the additional computation time and overhead introduced by non-end-to-end frameworks, as well as dependence on the capability of feature extractor. In contrast, our method is fully end-to-end, eliminating the need for additional training processes and completely removing the reliance on feature extractor initialization.
II-C Methodological Parallels from Self-Supervised ReID
ReID [41] aims to match the same person image captured by non-overlapping cameras. Due to its critical role in video surveillance, ReID has attracted extensive research attention and achieved remarkable progress. It has provided a fully end-to-end framework to eliminate the need for additional training processes and completely remove the reliance on feature extractor initialization, which offers strong inspiration for our consequent work in the DVGL task. However, these achievements are facilitated by extensive human-labeled data. To address this limitation, self-supervised learning methods [5, 6] have been proposed. Among them, clustering algorithms are widely employed to generate pseudo-labels, which serve as supervision signals for model training. In addition, benefitting from the introduction of instance-level contrastive learning [42] with memory banks and momentum update strategies [43], many methods [44, 45] consider each unlabeled sample as an individual class to learn discriminative instance-level feature representations. However, these instance-level features were largely modified during training, which resulted in unstable representations [46]. To address this issue, recent studies [47, 48] have begun exploring cluster-level relationships to discover more stable and robust associations among pedestrian representations. Cluster-Contrast [17] contrasts instance features with these cluster-level representations to address the issue of inconsistent instance discrimination. Nevertheless, in cross-modal ReID scenarios, where images are captured by heterogeneous sensors, modality discrepancies hinder the effectiveness of these methods in learning modality-invariant features. ADCA [18] introduced a dual-path contrastive framework to bridge modality gaps and enhance cross-modal feature alignment. Building on this, SDCL [49] employed a dual-path Transformer for shallow-to-deep contrastive learning, and refined pseudo-labels via consistency constraints across feature hierarchies to achieve more stable and discriminative cross-modal representations. The above methods focus on alleviating the dependency on human-labeled data in the ReID and provide valuable insights for DVGL. However, DVGL suffers from more severe viewpoint and scale variations. Furthermore, the spatial continuity of geographic environments inevitably induces feature ambiguity, which would severely degrade clustering effectiveness.
A cluster-based self-supervised framework, which is developed concurrently with our work, is proposed with a similar ReID framework that employs intermediate state transition and local-to-global matching for cross-view alignment to mitigate viewpoint discrepancies. However, this method relies on explicit appearance-level transformations rather than intrinsically modeling cross-view semantic features, and lacks end-to-end optimization and uncertainty-aware pseudo-label refinement, making it sensitive to noise and limiting its adaptability. In this work, we integrate instance-level and cluster-level features in a self-supervised method for DVGL by employing reliable neighbor learning to facilitate implicit cross-view alignment and improve the discriminative power and robustness of the learned representations.
Figure 3: The overall pipeline of our method consists of a lightweight backbone, a dual-path contrastive learning strategy, a dynamic hierarchical memory learning module, and an information consistency evolution module. Specifically, the dual-path contrastive learning is designed to learn discriminative and consistent intra-view feature representations. The dynamic hierarchical memory module further captures intra-view feature variations under different perspectives and scales, thereby enhancing the robustness and discriminability of the learned representations. The information consistency evolution module focuses on modeling cross-view feature consistency through a neighborhood-driven learning strategy, and further improves the training process by integrating a pseudo-label enhancement strategy.
III End-to-End Self-Supervised Learning Method
As illustrated in Fig. 3, the proposed method is composed of two key modules, DHML and ICEL. Additionally, a PLE strategy is employed to optimize the self-supervised learning process. Specifically, we adopt a weight-shared ConvNeXt-Tiny network as the backbone, equipped with cluster memory and a dual-path contrastive learning strategy to learn intra-view feature representations. To capture the dynamic variations of intra-class features caused by changes in viewpoint and scale, the DHML module establishes a collaborative relationship between short-term and long-term memory. DHML module enables efficient discriminability of intra-view feature representations. Furthermore, the ICEL module leverages instance memory to construct neighborhood-driven dynamic constraints. By integrating neighborhood consistency with mutual information optimization, it guides cross-view features to form consistent and unified representations.
Problem Formulation: In DVGL, a set of paired drone-satellite images is denoted as $\{x^{d},x^{s}\}$, where $x^{d}$ represents the drone-view image and $x^{s}$ represents the satellite-view image. Unlike the supervised setting, our self-supervised method is developed without using paired drone-satellite images, where the correspondences between drone-view and satellite-view images are entirely unavailable. The objective is to discover reliable cross-view correspondences under such conditions and to accomplish the DVGL task.
III-A Dual-Path Contrastive Learning Baseline
Cluster-based contrastive learning methods have achieved remarkable success in various fields [17, 50]. However, in DVGL, the unique challenges of cross-view scenarios, as discussed in Section I, make it difficult for single-path cluster-based contrastive learning to effectively capture view-specific features. Therefore, we design the baseline network with a dual-path architecture, where drone-view and satellite-view features are extracted using a shared backbone and clustered to assign pseudo-labels. Subsequently, a dual-path contrastive learning strategy [51] is employed to enhance intra-view representation learning.
We first introduce the notation for better clarity. Let $X_{d}=\{x^{d}_{i}\}_{i=1}^{N}$ denote a set of drone images with $N$ instances, and $X_{s}=\{x^{s}_{j}\}_{j=1}^{M}$ denote a set of satellite images with $M$ instances. $F_{d}=\{\mathbf{f}^{d}_{i}\}_{i=1}^{N}$ and $F_{s}=\{\mathbf{f}^{s}_{j}\}_{j=1}^{M}$ represent the feature representations extracted by the shared feature extractor $\mathcal{F}_{backbone}$ for drone-view and satellite-view, respectively. $q^{d}$ and $q^{s}$ denote the drone query instance features and satellite query instance features extracted by the feature extractor $\mathcal{F}_{backbone}$.
Arbitrary Sampling and Backbone-based Feature Extraction. To simulate real-world scenarios where drone and satellite images are not paired, we independently and arbitrarily sample drone and satellite images at each training epoch to construct unpaired batches. Each image is subsequently processed by a shared backbone network to extract features. The backbone processing procedure can be described as
$\begin{split}&\mathbf{f}^{d}_{i}=\mathcal{F}_{\text{backbone}}(x^{d}_{i}),\quad i=1,\dots,N\\ &\mathbf{f}^{s}_{j}=\mathcal{F}_{\text{backbone}}(x^{s}_{j}),\quad j=1,\dots,M\end{split}$ (1)
Cross-View Cluster Memory Initialization. At the beginning of each training epoch, the feature memories for both views are initialized into $K$ clusters and $L$ clusters, respectively. Specifically, the cluster centroids of the drone-view $\{\phi_{1}^{d},\cdots,\phi_{K}^{d}\}$ and the satellite-view $\{\phi_{1}^{s},\cdots,\phi_{L}^{s}\}$ are stored in the drone memory dictionary $\mathcal{M}_{d}$ and the satellite memory dictionary $\mathcal{M}_{s}$, respectively, for view-specific memory initialization. This process can be formalized as
$\mathcal{M}_{d}\ \leftarrow\{\phi_{1}^{d},\cdots,\phi_{K}^{d}\},\mathcal{M}_{s}\ \leftarrow\{\phi_{1}^{s},\cdots,\phi_{L}^{s}\}$ (2)
$\phi_{k}^{d}=\frac{1}{|\mathcal{H}_{k}^{d}|}\sum_{\mathbf{f}_{i}^{d}\in\mathcal{H}_{k}^{d}}\mathbf{f}_{i}^{d},\phi_{l}^{s}=\frac{1}{|\mathcal{H}_{l}^{s}|}\sum_{\mathbf{f}_{j}^{s}\in\mathcal{H}_{l}^{s}}\mathbf{f}_{j}^{s}$ (3)
where $\mathcal{H}_{k}^{d}$ denotes the $k$-th cluster set in drone-view, $k=1,\cdots,K$, and $\mathcal{H}_{l}^{s}$ denotes the $l$-th cluster set in satellite-view, $l=1,\cdots,L$. ||$\cdot$|| ⋅ | indicates the number of instances per cluster.
Cross-View Cluster Memory Updating. During the training phase, we randomly sample $P$ different localization points from the training set, with each point comprising a fixed number $Z$ of instances. This sampling strategy results in a minibatch containing a total of $P\times Z$ query images. Subsequently, the query features extracted from the minibatch are integrated into the cluster representations via a momentum-based update mechanism. The cluster centroids $\phi_{k}^{d}$ and $\phi_{l}^{s}$ are updated iteratively as
$\quad\phi_{k}^{d}\leftarrow\alpha\phi_{k}^{d}+(1-\alpha)q^{d},\quad\phi_{l}^{s}\leftarrow\alpha\phi_{l}^{s}+(1-\alpha)q^{s}$ (4)
where $\alpha$ is the momentum updating factor.
Cross-View Cluster Memory Contrastive Loss. Given drone and satellite query $q^{d}$ and $q^{s}$, we compute the contrastive loss for drone-view and satellite-view as
$\mathcal{L}^{d}_{c1}=-\log\frac{\exp(q^{d}\cdot\phi^{d}_{+}/\tau)}{\sum_{k=0}^{K}\exp(q^{d}\cdot\phi^{d}_{k}/\tau)}$ (5)
$\mathcal{L}^{s}_{c1}=-\log\frac{\exp(q^{s}\cdot\phi^{s}_{+}/\tau)}{\sum_{k=0}^{K}\exp(q^{s}\cdot\phi^{s}_{l}/\tau)}$ (6)
where $\tau$ denotes the temperature hyper-parameter. Cluster centroids $\phi^{d}_{k}$ and $\phi^{s}_{l}$ are utilized as feature vectors at the cluster level to compute the distances between the query instance and all clusters. Here, $\phi^{d}_{+}$ and $\phi^{s}_{+}$ are positive cluster features associated with $q^{d}$ and $q^{s}$. The final optimization for baseline is denoted as
$\mathcal{L}_{c1}=\mathcal{L}^{d}_{c1}+\mathcal{L}^{s}_{c1}$ (7)
III-B Dynamic Hierarchical Memory Learning
Geographic environments inherently exhibit spatial continuity, where neighboring regions usually present similar visual patterns, while distant regions show significant differences. However, such spatial characteristics often introduce ambiguity in intra-view feature representations. To address this issue, we propose a DHML module that jointly exploits long-term and short-term memory to capture spatial dynamics and enhance the discriminability of intra-view features.
Consistent with the baseline network, we construct dynamic hierarchical memory learning in a dual-path manner, using drone feature extraction as an illustrative example. To build long-term memory, we first construct two novel clustering memory dictionaries according to Eq.(2), which are used to store short-term $\mathcal{M}_{d}^{{}^{\prime}}$ and long-term $\mathcal{M}_{d}^{{}^{\prime\prime}}$ memories, respectively. They are expressed as
$\mathcal{M}_{d}^{{}^{\prime}}\ \leftarrow\{\varphi_{1}^{d{{}^{\prime}}},\cdots,\varphi_{K}^{d{{}^{\prime}}}\},\mathcal{M}_{d}^{{}^{\prime\prime}}\ \leftarrow\{\varphi_{1}^{d{{}^{\prime\prime}}},\cdots,\varphi_{K}^{d{{}^{\prime\prime}}}\}$ (8)
where $\varphi_{1}^{d{{}^{\prime}}},\dots,\varphi_{K}^{d{{}^{\prime}}}$ represent the short-term memory embeddings that encode the cluster representations of the drone-view, and $\varphi_{1}^{d{{}^{\prime\prime}}},\dots,\varphi_{K}^{d{{}^{\prime\prime}}}$ represent the long-term memory embeddings that store historically accumulated stable representations of the drone-view.
Unlike the weight updating scheme in Eq.(4), we assign different weights to the input features and historical memory during the long-term memory update. Specifically, since the momentum parameter is set to 0.2 throughout this study, we enhance the influence of historical memory during the initialization of long-term memory updates. The update procedure can defined as
$\quad\varphi_{k}^{d{{}^{\prime\prime}}}\leftarrow(0.5-\alpha)\varphi_{k}^{d{{}^{\prime\prime}}}+\alpha q^{d{{}^{\prime\prime}}}$ (9)
where $q^{d{{}^{\prime\prime}}}$ represents the drone query instance features. This obtains a new dictionary of long-term memory.
In addition, we further introduce an adaptive weight update mechanism to update the long-term memory. In detail, we compute the euclidean distance between the current input and its corresponding long-term memory, then take the average over all samples in the batch. This average distance is then mapped to the range $[0,1]$ via the sigmoid function, yielding an adaptive update coefficient:
$\beta=\delta\!\left(\frac{1}{B}\sum_{i=1}^{B}\|q^{d{{}^{\prime\prime}}}-\varphi_{k}^{d{{}^{\prime\prime}}}\|\right)$ (10)
where $B$ denotes the batch size and $\delta(\cdot)$ represents the sigmoid function. The short-term memory is subsequently updated as
$\quad\varphi_{k}^{d{{}^{\prime}}}\leftarrow\beta\cdot\varphi_{k}^{d{{}^{\prime\prime}}}+(1-\beta)\cdot\varphi_{k}^{d{{}^{\prime}}}$ (11)
Then, the long-term and short-term memories are fused in a weighted manner to yield the combined feature $\varphi_{k}^{db}$ representation:
$\varphi_{k}^{d{b}}=w{{}^{\prime\prime}}\cdot\varphi_{k}^{d{{}^{\prime\prime}}}+w{{}^{\prime}}\cdot\varphi_{k}^{d{{}^{\prime}}}$ (12)
where $w{{}^{\prime\prime}}$ and $w{{}^{\prime}}$ denote the respective fusion weights. As a result, we obtain a comprehensive memory dictionary that effectively integrates long-term global consistency with localized short-term adaptability.
Given the drone query $q^{d}$, we compute the contrastive loss by measuring their similarity to the cluster representations maintained in the memory dictionary by the following equations:
$\mathcal{L}^{d}_{c2}=-\log\frac{\exp(q^{d}\cdot\varphi^{db}_{+}/\tau)}{\sum_{k=0}^{K}\exp(q^{d}\cdot\varphi^{db}_{k}/\tau)}$ (13)
where $\tau$ denotes the temperature hyper-parameter. The combined cluster centroids $\varphi^{db}_{k}$ are used as feature vectors at the cluster level to calculate the distances between the query instance $q^{d}$ and all clusters. Here, $\varphi^{db}_{+}$ corresponds to the positive cluster feature associated with $q^{d}$. Finally, we introduce a fusion factor $\lambda_{0}$ to perform a weighted fusion of $\mathcal{L}^{d}_{c2}$ and $\mathcal{L}^{d}_{c1}$, aiming to enhance the discriminability of intra-view feature representations and obtain the optimized loss $\mathcal{L}_{c3}^{d}=\lambda_{0}\mathcal{L}^{d}_{c1}\ +\mathcal{L}^{d}_{c2}$. Similarly, $\mathcal{L}_{c3}^{s}$ can be obtained in the same way. Therefore, the dynamic hierarchical memory learning process can be summarized as
$\mathcal{L}_{c3}=\mathcal{L}^{d}_{c3}+\mathcal{L}^{s}_{c3}$ (14)
III-C Information Consistency Evolution Learning
The baseline network and the DHML module primarily focus on intra-view feature learning, lacking the ability to model and associate cross-view features effectively. Therefore, the proposed ICEL module employs a threshold-based and ranking-based neighborhood selection strategy to achieve effective cross-view feature alignment. By promoting the consistency evolution of cross-view feature representations, ICEL enhances both the discriminability and the representation capability of cross-view features.
In the ICEL module, unlike the baseline and DHML module, we employ instance-level features instead of cluster-level representations. To achieve this, we define two instance memory dictionaries, $\mathcal{I}_{d}$ and $\mathcal{I}_{s}$, which store the instance features extracted by the shared feature encoder, as described in Section III-A. This process can be described as
$\mathcal{I}_{d}\leftarrow\{\mathbf{f}_{1}^{d},\dots,\mathbf{f}_{N}^{d}\},\mathcal{I}_{s}\leftarrow\{\mathbf{f}_{1}^{s},\dots,\mathbf{f}_{M}^{s}\}$ (15)
where $\mathbf{f}_{i}^{d}$, $i=1,\dots N$, and $\mathbf{f}_{j}^{s}$, $j=1,\dots M$, denote the drone and satellite instance features stored in memory dictionaries $\mathcal{I}_{d}$ and $\mathcal{I}_{s}$, respectively.
To facilitate the following description, we take the scenario where the drone-view serves as the query as an illustrative example. The similarity between a given drone query $q_{v}^{d}$ that denotes the $v$-th query sample in the drone-view, $v=1,\dots N^{b}$. $N^{b}$ denote the batch size of the query $q_{d}$. Each drone instance $\mathbf{f}_{i}^{d}$ and each satellite instance $\mathbf{f}_{j}^{{s}}$ in the training set are compared to drone query $q_{d}$ by computing a similarity score, respectively. This is defined as
$\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d})=\frac{q_{v}^{d}\cdot\mathbf{f}_{i}^{d}}{\|q_{v}^{d}\|_{2}\|\mathbf{f}_{i}^{d}\|_{2}},\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s})=\frac{q_{v}^{d}\cdot\mathbf{f}_{j}^{s}}{\|q_{v}^{d}\|_{2}\|\mathbf{f}_{j}^{s}\|_{2}}$ (16)
Then, we adopt a threshold-based neighborhood selection strategy to select reliable neighbors, which can be defined as
$\Omega^{dd}=\Big{\{}\mathbf{f}_{i}^{d}\,\Big{|}\,S\big{(}q_{v}^{d},\mathbf{f}_{i}^{d}\big{)}>\gamma\max_{i\in\{1,\dots,N\}}S\big{(}q_{v}^{d},\mathbf{f}_{i}^{d})\Big{\}}$ (17)
$\Omega^{ds}=\Big{\{}\mathbf{f}_{j}^{s}\,\Big{|}\,S\big{(}q_{v}^{d},\mathbf{f}_{j}^{s}\big{)}>\gamma\max_{j\in\{1,\dots,M\}}S\big{(}q_{v}^{d},\mathbf{f}_{j}^{s})\Big{\}}$ (18)
where $\Omega^{dd}$ represents the set of drone feature instances selected as valid intra-view neighbor, and $\Omega^{ds}$ denotes the set of drone feature instances that exhibit strong cross-view associations with satellite features. $\gamma\in(0,1)$ is a predefined threshold.
Additionally, we introduce a ranking-based neighborhood selection strategy. Specifically, we independently select the top-$k_{1}$ most relevant samples to construct a strictly filtered neighborhood, ensuring high-confidence associations, while also selecting the top-$k_{2}$ samples define an expanded neighborhood, capturing additional informative relationships. The process can be defined as
$\mathcal{N}_{k_{1}}^{dd}=\Big{\{}\mathbf{f}_{i}^{d}\mid i\in\arg\max_{I\subset\{1,\dots,N\},|I|=k_{1}}\sum_{i\in I}S(q_{v}^{d},\mathbf{f}_{i}^{d})\Big{\}}$ (19)
$\mathcal{N}_{k_{1}}^{ds}=\Big{\{}\mathbf{f}_{j}^{s}\mid j\in\arg\max_{J\subset\{1,\dots,M\},|J|=k_{1}}\sum_{j\in J}S(q_{v}^{d},\mathbf{f}_{j}^{s})\Big{\}}$ (20)
where $\mathcal{N}_{k_{1}}^{dd}$ and $\mathcal{N}_{k_{1}}^{ds}$ denote the selected intra-view and cross-view neighborhood sets, respectively. $\mathcal{N}_{k_{2}}^{dd}$ and $\mathcal{N}_{k_{2}}^{ds}$ can be obtained following the same selection strategy.
Based on the constructed neighborhoods, the normalized similarity distribution is computed within the larger neighborhood to promote the aggregation of similar samples in both intra-view and cross-view scenarios. This distribution is then optimized by a cross-entropy loss to enhance alignment between query samples and their neighbors. The intra-view $\mathcal{L}_{\Omega}^{dd}$ and cross-view $\mathcal{L}_{\Omega}^{ds}$ losses can be formulated as
$\mathcal{L}_{\Omega}^{dd}\!=\!-\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{\mathbf{f}_{i}^{d}\in\Omega^{dd}}\!\log\frac{\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{i}^{d})/\tau\right)}{\sum_{\mathbf{f}_{i}^{d}\in{{{\Omega}^{dd}}}}\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{i}^{d})/\tau\right)}$ (21)
$\mathcal{L}_{\Omega}^{ds}\!=\!-\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{\mathbf{f}_{j}^{s}\in\Omega^{ds}}\!\log\frac{\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{j}^{s})/\tau\right)}{\sum_{\mathbf{f}_{j}^{s}\in{{{\Omega}^{ds}}}}\exp\left(\mathcal{S}(q_{v}^{d{}},\mathbf{f}_{j}^{s})/\tau\right)}$ (22)
where $N^{{\Omega}^{dd}}$ and $N^{{\Omega}^{ds}}$ denote the total number of neighborhood samples within the set ${\Omega}^{dd}$ and ${\Omega}^{ds}$.
Incorporating neighborhoods $\mathcal{N}_{k_{1}}^{dd}$, $\mathcal{N}_{k_{2}}^{dd}$ and $\mathcal{N}_{k_{1}}^{ds}$, $\mathcal{N}_{k_{2}}^{ds}$, we introduce consistency loss to enhance alignment of relative similarity distributions within the large-scale neighborhood $k_{2}$ across both intra-view and cross-view representations. The intra-view loss $\mathcal{L}_{k_{2}}^{dd}$ and cross-view loss $\mathcal{L}_{k_{2}}^{ds}$ under this consistency loss can be formulated as
$p_{v,i}^{dd}=\frac{\exp(\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d}))}{\sum_{\mathbf{f}_{i}^{d}\in\mathcal{N}_{k_{2}}^{dd}}\exp(S(q_{v}^{d},\mathbf{f}_{i}^{d}))}$ (23)
$p_{v,j}^{ds}=\frac{\exp(\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s}))}{\sum_{\mathbf{f}_{j}^{s}\in\mathcal{N}_{k_{2}}^{ds}}\exp(S(q_{v}^{d},\mathbf{f}_{j}^{s}))}$ (24)
$\mathcal{L}_{k_{2}}^{dd}=\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{i\in\mathcal{N}_{k_{2}}^{dd}}p_{v,i}^{dd}\log\left(k_{2}\cdot p_{v,i}^{dd}\right)$ (25)
$\mathcal{L}_{k_{2}}^{ds}=\frac{1}{N^{b}}\sum_{v=1}^{N^{b}}\sum_{j\in\mathcal{N}_{k_{2}}^{ds}}p_{v,j}^{ds}\log\left(k_{2}\cdot p_{v,j}^{ds}\right)$ (26)
where $p_{v,i}^{dd}$ and $p_{v,j}^{ds}$ are the softmax-normalized similarity distribution.
Within a more restrictive and smaller set of $k_{1}$ neighbors, a mutual information constraint is proposed to learn intra-view and cross-view features representations. This complements the consistency constraint described above and achieves a balance between global structural alignment and local discriminative information. Within the selected $k_{1}$ neighbors, we compute the similarity between the query $q_{v}^{d}$ and each neighbor $\mathbf{f}_{i}^{d}$ and $\mathbf{f}_{j}^{s}$ using the function $S(\cdot,\cdot)$, and apply softmax to obtain the posterior probability for each neighbor. It can be described as
$p^{dd}(i|v)=\frac{\exp\big{(}\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d})\big{)}}{\sum_{\mathbf{f}_{i}^{d}\in\mathcal{N}_{k_{1}}^{dd}}\exp\big{(}\mathcal{S}(q_{v}^{d},\mathbf{f}_{i}^{d})\big{)}}$ (27)
$p^{ds}(j|v)=\frac{\exp\big{(}\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s})\big{)}}{\sum_{\mathbf{f}_{j}^{s}\in\mathcal{N}_{k_{1}}^{ds}}\exp\big{(}\mathcal{S}(q_{v}^{d},\mathbf{f}_{j}^{s})\big{)}}$ (28)
Therefore, the intra-view and cross-view mutual information losses $\mathcal{L}_{k_{1}}^{dd}$ and $\mathcal{L}_{k_{1}}^{ds}$ can be derived and are defined as
$\mathcal{L}_{k_{1}}^{dd}=-\frac{1}{N_{b}}\sum_{v=1}^{N_{b}}\sum_{i=1}^{k_{1}}p^{dd}(i|v)\log\frac{p^{dd}(i|v)}{1/R}$ (29)
$\mathcal{L}_{k_{1}}^{ds}=-\frac{1}{N_{b}}\sum_{v=1}^{N_{b}}\sum_{j=1}^{k_{1}}p^{ds}(j|v)\log\frac{p^{ds}(j|v)}{1/R}$ (30)
For each view, the loss is formulated as a weighted sum of the original neighborhood loss $\mathcal{L}_{\Omega}$, the strict neighborhood loss $\mathcal{L}_{k_{1}}$, and the extended neighborhood loss $\mathcal{L}_{k_{2}}$. Specifically, the intra-view loss $\mathcal{L}^{dd}$ and the cross-view loss $\mathcal{L}^{ds}$ are respectively defined as
$\mathcal{L}_{c4}^{dd}=\mathcal{L}_{\Omega}^{dd}+\lambda_{1}\mathcal{L}_{k_{1}}^{dd}+\lambda_{2}\mathcal{L}_{k_{2}}^{dd}$ (31)
$\mathcal{L}_{c4}^{ds}=\mathcal{L}_{\Omega}^{ds}+\lambda_{1}\mathcal{L}_{k1}^{ds}+\lambda_{2}\mathcal{L}_{k2}^{ds}$ (32)
where $\lambda_{1}$ and $\lambda_{2}$ are hyperparameters controlling the contribution of strict and extended neighborhood components. When the satellate view as the query, the $\mathcal{L}_{c4}^{sd}$ and $\mathcal{L}_{c4}^{ss}$ can be obtained in the same way. The total neighbor learning loss is computed as $\mathcal{L}_{c4}=\mathcal{L}_{c4}^{dd}+\mathcal{L}_{c4}^{ds}+\mathcal{L}_{c4}^{sd}+\mathcal{L}_{c4}^{ss}$.
Finally, the overall loss of our method is defined as
$\mathcal{L}_{total}=\mathcal{L}_{c1}+\mathcal{L}_{c3}+\mathcal{L}_{c4}$ (33)
III-D Pseudo-Label Enhancement
In self-supervised learning, the quality of pseudo-labels exerts a decisive influence on model performance. Inspired by [49], we propose a pseudo-label optimization strategy based on dynamic feature robustness. Specifically, we introduce small perturbations to the original feature representations and exploit the top-K neighbor consistency between the original and perturbed features to identify high-confidence samples while filtering out noisy instances. Then, a spatial smoothing operation based on the similarity matrix is applied to refine the pseudo-labels to enhance local label consistency.
Small Gaussian noise $\boldsymbol{\epsilon}_{d},\boldsymbol{\epsilon}_{s}\sim\mathcal{N}(\mathbf{0},\sigma^{2}\mathbf{I})$ is used in the initial satellite instances feature and the initial drone instances features to obtain perturbed features, where the $\sigma$ denotes the standard deviation. It can be formulated as
$\mathbf{f}_{i}^{d{\boldsymbol{\epsilon}}}=\mathbf{f}_{i}^{d}+\boldsymbol{\epsilon}_{d},\quad\mathbf{f}_{j}^{s{\boldsymbol{\epsilon}}}=\mathbf{f}_{j}^{s}+\boldsymbol{\epsilon}_{s}$ (34)
Using the definition of the similarity score shown in Eq.(16), two cross-view ranking lists can be obtained, which are denoted as $\mathcal{K}^{ds}=\{\mathcal{S}(\mathbf{f}_{i}^{d},\mathbf{f}_{j}^{s})\}$ and $\mathcal{K}^{ds}_{\boldsymbol{\epsilon}}=\{\mathcal{S}(\mathbf{f}_{i}^{d{\boldsymbol{\epsilon}}},\mathbf{f}_{j}^{s{\boldsymbol{\epsilon}}}))\}$, $i=1,\cdots,N$. The label of the $k$-th similar drone instance in two ranking lists can be denoted as $y_{\mathbf{f}_{i}^{d}}[k]$ and $y_{\mathbf{f}_{i}^{d{\boldsymbol{\epsilon}}}}[k]$. Next, the cross-view labels are associated with the intersection of two label sets to investigate collaborative ranking consistency. This process can be described as
$I_{\mathbf{f}_{i}^{d}}=\{y_{\mathbf{f}_{i}^{d}}[k]\cap y_{\mathbf{f}_{i}^{d{\boldsymbol{\epsilon}}}}[k]\}_{k=1}^{N}$ (35)
where $I_{\mathbf{f}_{i}^{d}}$ records the samples with the same identity in instances. Accordingly, we update the cross-view pseudo-label for $\mathbf{f}_{j}^{s}$ using the label with the highest frequency in $I_{\mathbf{f}_{i}^{d}}$, which reflects the most stable identity under both original and perturbed feature rankings. This final refined label is denoted as $\tilde{y}_{\mathbf{f}_{j}^{s}}$. Then, we convert the label list $\{\tilde{y}_{\mathbf{f}_{j}^{s}}\}$ into the form of one-hot code matrix $\tilde{\mathbf{y}}\in\mathbb{R}^{M\times N}$ by setting the column according to the refined labels to 1 and the rest to 0.
In order to explore the intra-view initial and perturbed collaboration, the initial and perturbed similarity matrices $\mathbf{P}^{ss}$ and $\mathbf{P}^{ss}_{\boldsymbol{\epsilon}}$ are calculated, and their sum $\mathbf{P}$ is used to investigate intra-view ranking consistency. Then we keep the 5-max values of $\mathbf{P}$ in each row at 1 and set the rest to 0, acquiring the ranking relations. The process of intra-view ranking smoothing is formulated as
$\mathbf{y}=\mathbf{P}\tilde{\mathbf{y}}$ (36)
where $\mathbf{y}\in\mathbb{R}^{M\times N}$ is the final refined cross-view label matrix of satellite instance. In $\mathbf{y}$, the column number of the maximum value in each row is the refined label of the samples.
TABLE I: Comparisons between the proposed method and some state-of-the-art methods on the University-1652 datasets.
University-1652
[…]
Model Venue Supervision Ratio Parameters(M) R@1 AP R@1 AP
MuSe-Net[52] PR’2024 100(%) 50.47 74.48 77.83 88.02 75.10
MCCG[1] TCSVT’2023 100(%) 56.65 89.40 91.07 95.01 89.93
SDPL[53] TCSVT’2024 100(%) 42.56 90.16 91.64 93.58 89.45
CCR[31] TCSVT’2024 100(%) 156.57 92.54 93.78 95.15 91.80
Sample4Geo[28] ICCV’2023 100(%) 87.57 92.65 93.81 95.14 91.39
SRLN[30] TGRS’2024 100(%) 193.03 92.70 93.77 95.14 91.97
MEAN[2] arXiv’2024 100(%) 36.50 93.55 94.53 96.01 92.08
DAC[3] TCSVT’2024 100(%) 96.50 94.67 95.50 96.43 93.79
Ours - 100(%) 28.59 92.02 93.36 95.57 91.27
EM-CVGL[4] TGRS’2024 0(%) 2.25 70.29 74.93 79.03 61.03
CDIKTNet[40] arXiv’2025 0(%) 91.52 83.30 85.73 87.73 76.53
Wang[9] AAAI’2025 0(%) - 85.95 90.33 94.01 82.66
Ours - 0(%) 28.59 90.17 91.67 95.01 88.95
IV Experimental results
IV-A Experimental Datasets and Evaluation Metrics
To evaluate the proposed method, we conduct experiments on three benchmark datasets, including University-1652 [12], SUES-200 [32] and DenseUAV [52].
University-1652 contains abundant images of 1,652 buildings in 72 universities around the world. It involves drone, satellite, and ground-level views. The training set consists of 701 buildings from 33 universities and the test set covers 951 buildings from 39 geographically nonaffiliated universities.
SUES-200 focuses on altitude diversity in aerial imagery, containing 200 locations, with 120 designated for training and 80 for testing. Each location in the dataset contributes one satellite image along with aerial images captured at four distinct altitudes, which involve urban areas, natural landscapes, and water bodies. SUES-200 evaluates cross-view retrieval under varying flight heights, reflecting real-world drone scenarios.
DenseUAV focuses on drone self-positioning in low-altitude urban environments with more than 27,000 drone and satellite images collected from 14 university campuses. It is characterized by dense sampling, multi-scale satellite imagery, and diverse altitudes captured across multiple temporal conditions. DenseUAV poses challenges such as cross-view matching, neighbor confusion, and spatial-temporal variation.
We evaluate the models using Recall@K (R@K) and Average Precision (AP), where R@K measures the proportion of correct matches among the Top-$K$results and AP evaluates the overall precision-recall performance. Model efficiency is assessed by the number of parameters to reflect portability under resource constraints. All comparisons are conducted using the best configurations for each method.
IV-B Implementation Details
A ConvNeXt-Tiny model, pre-trained on ImageNet, serves as the backbone for feature extraction, with a newly added classifier module initialized via the Kaiming method. During both training and testing, all input images are resized to $3\times 384\times 384$. We employ several data-augmentation techniques, including random cropping, random horizontal flipping, and random rotation. The batch size is set to 64, and DBSCAN [7] is used to generate pseudo-labels. For optimization, we use the SGD optimizer with an initial learning rate of 0.001, training the model for 30 epochs in total. All experiments are conducted under the PyTorch framework on an Ubuntu 22.04 system equipped with four NVIDIA RTX 4090 GPUs. Additionally, for the University-1652 and SUES-200 datasets, we apply feature replication only on the training sets by duplicating each satellite feature 50 times for clustering purposes, while drone features remain unchanged. The test sets are kept intact without any modification. In the DenseUAV dataset, drone features are replicated four times at each altitude, and satellite features are replicated three times for each time point and altitude exclusively in the training set; the test set remains unchanged.
IV-C Comparison with State-of-the-art Methods
To evaluate the effectiveness of our method, we compare it with the state-of-the-art supervised methods MuSe-Net [52], MCCG [1], SDPL [53], CCR [31], Sample4Geo [28], SRLN [30], MEAN [2] and DAC [3], and the self-supervised methods EM-CVGL [4], CDIKTNet [40] and Wang [9].
[…]
We further benchmark with eight representative supervised methods, and the results show that our method clearly surpasses some of them, including MuSe-Net, MCCG, and SDPL. Furthermore, with self-supervised pre-training and subsequent supervised fine-tuning, our method achieves competitive performance under the supervised setting, comparable to CCR, Sample4Geo, and SRLN.
[…]
SUES-200
[…]
150m 200m 250m 300m
Model Venue Supervision Ratio Parameters(M) R@1 AP R@1 AP R@1 AP R@1 AP
MCCG[1] TCSVT’2023 100(%) 56.65 82.22 85.47 89.38 91.41 93.82 95.04 95.07 96.20
SDPL[53] TCSVT’2024 100(%) 42.56 82.95 85.82 92.73 94.07 96.05 96.69 97.83 98.05
CCR[31] TCSVT’2024 100(%) 156.57 87.08 89.55 93.57 94.90 95.42 96.28 96.82 97.39
SRLN[30] TGRS’2024 100(%) 193.03 89.90 91.90 94.32 95.65 95.92 96.79 96.37 97.21
Sample4Geo[28] ICCV’2023 100(%) 87.57 92.60 94.00 97.38 97.81 98.28 98.64 99.18 99.36
DAC[3] TCSVT’2024 100(%) 96.50 96.80 97.54 97.48 97.97 98.20 98.62 97.58 98.14
MEAN[2] arXiv’2024 100(%) 36.50 95.50 96.46 98.38 98.72 98.95 99.17 99.52 99.63
Ours - 100(%) 28.59 94.00 95.21 97.83 98.88 99.14 95.44 99.25 99.39
EM-CVGL[4] TGRS’2024 0(%) 2.25 55.23 60.80 60.95 61.03 68.10 72.62 74.42 78.20
CDIKTNet[40] arXiv’2025 0(%) 91.52 82.75 85.25 89.35 91.13 93.15 94.39 95.18 96.12
Wang[9] AAAI’2025 0(%) - 76.90 84.95 87.88 92.60 92.98 95.66 95.10 96.92
Ours - 0(%) 28.59 82.60 85.60 90.48 92.28 94.33 95.44 96.78 97.42
Ours - 0(%) 28.59 87.00 89.20 93.98 95.11 96.50 97.23 97.23 97.86
Ours - 0(%) 28.59 90.98 92.66 95.03 96.14 97.45 98.03 98.23 98.64
[…]
150m 200m 250m 300m
Model Venue Supervision Ratio Parameters(M) R@1 AP R@1 AP R@1 AP R@1 AP
CCR[31] TCSVT’2024 100(%) 156.57 92.50 88.54 97.50 95.22 97.50 97.10 97.50 97.49
Sample4Geo[28] ICCV’2023 100(%) 87.57 92.60 94.00 97.38 97.81 98.28 98.64 99.18 99.36
MCCG[1] TCSVT’2023 100(%) 56.65 93.75 89.72 93.75 92.21 96.25 96.14 98.75 96.64
SDPL[53] TCSVT’2024 100(%) 42.56 93.75 83.75 96.25 92.42 97.50 95.65 96.25 96.17
SRLN[30] TGRS’2024 100(%) 193.03 93.75 93.01 97.50 95.08 97.50 96.52 97.50 96.71
DAC[3] TCSVT’2024 100(%) 96.50 97.50 94.06 98.75 96.66 98.75 98.09 98.75 97.87
MEAN[2] arXiv’2024 100(%) 36.50 97.50 94.75 100.00 97.09 100.00 98.28 100.00 99.21
Ours - 100(%) 28.59 97.50 92.71 97.50 97.43 100.00 98.98 98.75 98.25
EM-CVGL[4] TGRS’2024 0(%) 2.25 73.75 54.00 91.25 65.65 96.25 72.02 97.50 74.74
CDIKTNet[40] arXiv’2025 0(%) 91.52 88.75 80.33 93.75 88.17 95.00 92.15 98.75 94.37
Wang[9] AAAI’2025 0(%) - 87.50 74.81 92.50 87.15 96.25 91.20 98.75 94.52
Ours - 0(%) 28.59 91.25 81.28 97.50 90.62 98.75 94.74 97.50 96.17
Ours - 0(%) 28.59 95.00 85.46 97.50 92.72 100.00 95.43 100.00 96.39
Ours - 0(%) 28.59 96.25 90.05 97.50 94.80 98.75 96.82 98.75 97.33
Results on SUES-200. As shown in Table II, our method, when directly transferring the pre-trained model from the University-1652 dataset to the SUES-200 dataset without any re-training, consistently outperforms all existing self-supervised methods and achieves competitive performance compared to supervised methods such as MCCG and SDPL. Moreover, by performing self-supervised re-training on the SUES-200 dataset, our method brings a significant performance boost, achieving results comparable to the supervised CCR. Notably, when further fine-tuning the transferred model for only one epoch, we observe additional improvements, surpassing the supervised SRLN. These results clearly demonstrate the effectiveness of our method in self-supervised learning, transfer learning, and efficient fine-tuning.
Results on Cross-Domain Generalization Performance. As shown in Table III, our method consistently outperforms existing self-supervised methods in cross-domain settings, showing both significant improvements and stable generalization. Specifically, in the Drone to Satellite setting, our model achieves better performance than EM-CVGL at all altitudes. At 150 meters, R@1 and AP increase by approximately 22% and 20%. At 200 meters, the gains are around 19% and 16%; at 250 meters, nearly 14% and 12%. Even at 300 meters, our method still outperforms by about 11% and 9%. In the Satellite to Drone scenario, similar improvements are observed. At 150 meters, R@1 and AP improve by approximately 18% and 25%; at 200 meters, by 10% and 20%; at 250 meters, by 6% and 13%; and at 300 meters, by 3% and 10%. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes. Remarkably, our method even surpasses all existing state-of-the-art supervised methods in terms of cross-domain generalization. These results confirm the robustness and strong cross-domain generalization capability of our method across varying altitudes.
TABLE III: Comparisons between the proposed method and state-of-the-art methods in cross-domain evaluation.
[…]
150m 200m 250m 300m
Model Venue Supervision Ratio Parameters(M) R@1 AP R@1 AP R@1 AP R@1 AP
MCCG[1] TCSVT’2023 100(%) 56.65 57.62 62.80 66.83 71.60 74.25 78.35 82.55 85.27
Sample4Geo[28] ICCV’2023 100(%) 87.57 70.05 74.93 80.68 83.90 87.35 89.72 90.03 91.91
DAC[3] TCSVT’2024 100(%) 96.50 76.65 80.56 86.45 89.00 92.95 94.18 94.53 95.45
CAMP[29] TGRS’2024 100(%) 91.40 78.90 82.38 86.83 89.28 91.95 93.63 95.68 96.65
MEAN[2] arXiv’2024 100(%) 36.50 81.73 87.72 89.05 91.00 92.13 93.60 94.63 95.76
EM-CVGL[4] TGRS’2024 0(%) 2.25 60.03 65.69 71.50 76.17 80.35 83.85 85.93 88.34
Ours - 0(%) 28.59 82.60 85.60 90.48 92.28 94.33 95.44 96.78 97.42
[…]
150m 200m 250m 300m
Model Venue Supervision Ratio Parameters(M) R@1 AP R@1 AP R@1 AP R@1 AP
MCCG[1] TCSVT’2023 100(%) 56.65 61.25 53.51 82.50 67.06 81.25 74.99 87.50 80.20
Sample4Geo[28] ICCV’2023 100(%) 87.57 83.75 73.83 91.25 83.42 93.75 89.07 93.75 90.66
DAC[3] TCSVT’2024 100(%) 96.50 87.50 79.87 96.25 88.98 95.00 92.81 96.25 94.00
CAMP[29] TGRS’2024 100(%) 91.40 87.50 78.98 95.00 87.05 95.00 91.05 96.25 93.44
MEAN[2] arXiv’2024 100(%) 36.50 91.25 81.50 96.25 89.55 95.00 92.36 96.25 94.32
EM-CVGL[4] TGRS’2024 0(%) 2.25 73.75 56.99 87.50 70.62 92.50 81.18 95.00 86.04
Ours - 0(%) 28.59 91.25 81.28 97.50 90.62 98.75 94.74 97.50 96.17
TABLE IV: Comparisons between the proposed method and some state-of-the-art methods on the DenseUAV datasets.
DenseUAV
All height 80m 90m 100m
Model Venue Supervision Ratio Parameters(M) R@1 R@5 R@1 R@5 R@1 R@5 R@1 R@5
ConvNext-Tiny[33] TIP’2024 100(%) 30.10 60.23 81.94 - - - - - -
Sample4Geo[28] ICCV’2023 100(%) 87.57 80.57 96.53 80.57 96.53 86.87 98.71 91.38 99.74
CAMP[29] TGRS’2024 100(%) 91.40 88.72 97.76 76.45 96.01 84.81 99.23 89.19 99.23
DAC[3] TCSVT’2024 100(%) 96.50 84.47 96.53 80.92 97.04 85.32 98.71 85.46 99.10
MEAN[2] arXiv’2024 100(%) 36.50 90.18 97.86 78.76 97.04 85.33 99.10 89.32 98.97
Ours - 100(%) 28.59 86.31 97.81 86.10 98.46 88.93 99.61 90.35 99.49
EM-CVGL[4] TGRS’2024 0(%) 2.25 18.15 50.97 18.02 50.36 17.76 50.19 17.50 49.98
CDIKTNet[40] arXiv’2025 0(%) 91.52 22.57 40.24 17.25 40.15 23.04 48.39 25.48 52.38
Ours - 0(%) 28.59 73.79 90.69 73.10 92.92 74.65 95.37 75.03 95.50
Results on DenseUAV. The DenseUAV dataset presents a highly challenging scenario due to its densely sampled nature, where neighboring images exhibit high visual similarity, often leading to severe confusion in DVGL. In addition, the dataset includes satellite images captured at different times and drone images from multiple altitudes, which further increases intra-domain variability and task complexity. As shown in Table IV, in the fully unsupervised setting, our method achieves 73.79% R@1 and 90.69% R@5 across all altitudes, significantly outperforming existing self-supervised methods. It is noteworthy that the two representative self-supervised baselines consistently yield poor performance on DenseUAV, due to their limitations in handling densely sampled and spatio-temporal diverse scenarios. Furthermore, after supervised fine-tuning, our method surpasses several supervised baselines, including ConvNeXt-Tiny, Sample4Geo, CAMP, and DAC. In some cases, it even outperforms the current best-performing supervised method, MEAN. These results clearly demonstrate the effectiveness of our method in addressing extreme feature confusion caused by spatial continuity and appearance shifts induced by temporal variation, confirming its strong generalization and robustness.
IV-D Ablation Studies
To thoroughly evaluate the effectiveness of our proposed method, we perform an ablation study on the University-1652 dataset. We systematically assess the impact of each individual module and investigate their contributions under different configurations. The results are presented in Table LABEL:Ablation_Studies.
ConvNeXt-Tiny Backbone. The backbone refers to the untrained ConvNeXt-Tiny network used as a direct feature extractor. As shown in Table LABEL:Ablation_Studies, its performance is significantly limited, and obtains an R@1 of only 9.55% and 35.24%, and an AP of 11.88% and 13.20% for the Drone $\rightarrow$ Satellite and Satellite $\rightarrow$ Drone, respectively. These results indicate that the backbone alone lacks the ability to extract discriminative cross-view features.
Effectiveness of Dual-Path Contrastive Learning Baseline. After incorporating the baseline, the R@1 improves to 71.06% and 81.31%, while AP reaches 74.91% and 65.26%, respectively. Although the baseline primarily learns intra-view feature discrimination, the training process also facilitates a certain degree of alignment between drone and satellite representations. This leads to the formation of a more consistent and unified feature representation space, which serves as a solid foundation for subsequent cross-view learning.
Effectiveness of Dynamic Hierarchical Memory Learning. Compared with the baseline, the incorporation of DHML leads to significant performance improvements in both settings. Specifically, R@1 increases by 12.97% and 9.13%, and AP improves by 11.5% and 17.48% , reaching 84.03% and 90.44% in R@1, and 86.41% and 82.74% in AP, respectively. These gains can be attributed to the joint modeling of long-term and short-term feature dependencies introduced by DHML, which enhances the capacity to capture spatial structures in DVGL scenarios and enhances intra-view feature representation and discrimination.
TABLE V: The influence of each component on the performance of proposed method.
Setting University-1652
[…]
Backbone Baseline DHML ICEL PLE R@1 AP R@1 AP
9.55 11.88 35.24 13.20
71.06 74.91 81.31 65.26
84.03 86.41 90.44 82.74
74.48 77.96 89.87 74.37
86.13 88.31 94.01 82.97
90.17 91.67 95.01 88.95
Effectiveness of Information Consistency Evolution Learning. Although DHML enhances intra-view feature representation and discrimination, its optimization remains limited to intra-view information. To address this limitation, ICEL is proposed to learn cross-view dependencies. By enabling bidirectional information interaction between different views, ICEL facilitates cross-view feature refinement. As a result, R@1 improves to 86.13% and 51.04%, while AP reaches 88.31% and 82.97%. These results validate the effectiveness of ICEL in enhancing cross-view consistency for DVGL.
Pseudo-Label Enhancement Refining Feature Learning with Robust Labels. To mitigate pseudo-label noise caused by clustering, we apply PLE to improve label reliability. After incorporating PLE, our final model achieves state-of-the-art performance, with R@1 increasing to 90.17% and 95.01%, and AP reaching 91.67% and 88.95%. This confirms that refining pseudo-label quality significantly enhances the final model’s effectiveness in cross-geographic view matching.
The ablation study systematically evaluates the contribution of each component in our method. Starting with a backbone that lacks sufficient discriminative power, we first establish a dual-branch contrastive learning baseline. On top of this, DHML is introduced to strengthen intra-view feature representation and discriminability, ICEL is employed to facilitate cross-view feature association, and PLE is leveraged for pseudo-label refinement. Experimental results show that our method enhances the learning of discriminative and view-consistent representations under self-supervised learning.
Figure 4: The first row presents the t-SNE[54] visualization, while the second row illustrates the similarity distribution for randomly selected geographic locations. In the t-SNE visualization, each color represents a distinct location, where ’×’ markers correspond to the drone viewpoint, and the triangles indicate the satellite perspective of the same location.
IV-E Visualization
Feature space and similarity distribution. To further illustrate the effectiveness of the method, we perform a feature space and similarity distribution visualization for our method, as presented in Fig. 4. Comparing (a) Baseline and (b) Ours, it is evident that our method enhances the clustering of cross-view positive samples, effectively bringing the drone and satellite representations of the same location closer together. Additionally, the separation between different locations is more distinct that indicates better feature discriminability. The second row shows the similarity distribution between drone-satellite positive and negative pairs. Our method demonstrates a clearer distinction between positive and negative pairs, with a reduced overlap between distributions, which suggests improved cross-view matching and reduced feature ambiguity. This highlights the effectiveness of our method.
Figure 5: Cluster analysis across different epochs for various datasets. The plot shows the number of clusters for DenseUAV, University-1652, and SUES-200 datasets.
Cluster analysis. To validate the efficacy of our method, as shown in Fig. 5, we illustrate the evolution of cluster counts in each epoch for three datasets. In the DenseUAV dataset (2,255 locations), satellite images exhibit significant clustering fluctuations during the early training stages, due to low intra-class discriminability and weak cross-view correlations. In contrast, the University-1652 dataset (701 locations) demonstrates greater drone image clustering instability, while satellite image clusters stabilize more rapidly in the initial epochs. The SUES-200 dataset (120 locations), with its smaller data size and simpler distribution, exhibits minimal early-stage variation and stabilizes faster than larger datasets. As training progresses, the cluster counts across all datasets gradually converge toward the ground-truth location counts, despite initial inter-class confusion. This process validates the ability of our method to learn discriminative representations progressively.
Figure 6: Top-6 retrieval results of the proposed method on the University-1652 dataset.
[…]
Compared with University-1652, the DenseUAV dataset introduces more realistic and complex challenges, such as dense spatial sampling, multi-scale resolutions, and temporal dynamics. In the drone $\rightarrow$ satellite retrieval setting, despite the high risk of confusion caused by dense spatial sampling, our method effectively captures spatial-geographic cues and achieves high retrieval accuracy, demonstrating robust cross-view performance under realistic and complex conditions.
Figure 7: Top-6 retrieval results of the proposed method on the DenseUAV dataset.
V CONCLUSION
This paper investigates a highly valuable and challenging DVGL task to reduce reliance on paired drone–satellite images. To this end, we propose a self-supervised learning method that employs a shallow backbone with only 9 layers for feature extraction. Pseudo-labels are generated via clustering, and a dual-branch contrastive learning mechanism is introduced to enhance the capability in representing intra-view visual patterns. To further improve intra-view discriminability and consistency, we design a dynamic hierarchical memory learning module. Additionally, to mitigate the structural and semantic domain gap between views, we propose an information consistency evolution learning module, which facilitates semantic alignment by learning high-dimensional feature consistency across-views. To improve the quality and robustness of pseudo-labels, we introduce a pseudo-label enhancement strategy, which refines pseudo-supervision signals and boosts cross-view representation learning. As a result, the proposed method effectively unifies both intra-view and cross-view feature representations in the embedding space. Extensive experiments conducted on three public benchmarks University-1652, SUES-200, and DenseUAV, demonstrate that our method significantly outperforms existing self-supervised methods, and even surpasses several fully supervised methods in certain scenarios without relying on paired drone–satellite images. These results validate the effectiveness and generalization capability of our method and verify the potential for real-world DVGL applications.